solve by the elimination method.
5x+2y=-11
7x-3y=17

Found 2 solutions by richwmiller, stanbon:
Answer by richwmiller(17219)

5x+2y=-11
7x-3y=17
To solve by the elimination method (not the substitution method) we add, subtract multiply or divide one whole equation to remove either x or y (eliminate x or y)
I suggest multiplying the first equation by 3 and the second by 2 and adding them
5x+2y=-11
by 3
15x+6y=-33
7x-3y=17
15x+6y=-33
14x-6y=34
add them
29x=1
x=1/29
plug that in

Answer by stanbon(75887)

solve by the elimination method.
5x+2y=-11
7x-3y=17
--------------
Multiply thru 1st by 3 and thru 2nd by 2:
15x + 6y = -33
14x - 6y = 34
--------------------
29x = 1
x = 1/29
--------------
Substitute into 5x+2y = -11 to solve for "y":
5(1/29) + 2y = -11
Multiply thru by 29 to get rid of the denominator:
5 + 58y = -319
58y = -324
y = -162/29
=======================
Cheers,
Stan H.